7. Two random samples drawn from two normal populations are
Sample I 20 16 26 27 23 22 18 24 25 19 - -
Sample II 27 33 42 35 32 34 38 28 41 43 30 37
Obtain the estimates of the variances of the population.
1
Expert's answer
2020-05-14T17:46:17-0400

"S^{2}=\frac{1}{n-1}(\sum x^{2}-\frac{(\sum x)^{2}}{n})\\
\text{For the first sample n=10}\\
S_{1}^{2}=\frac{1}{9}(20^{2}+16^{2}+...+19^{2}-\frac{(20+16+...+19)^{2}}{10})\\
=13.33\\
\text{For the second sample n=12}\\
S_{2}^{2}=\frac{1}{11}(27^{2}+33^{2}+...+37^{2}-\frac{(27+33+...+37)^{2}}{12})\\
=28.55\\
S_{p}^{2}=\frac{(n_{1}-1)S_{1}^{2}+(n_{2}-1)S_{2}^{2}}{(n_{1}+n_{2}-2)}\\
S_{p}^{2}=\frac{(9)(13.33)+(11)(28.55)}{20}\\
=21.701"